# RLGNet: Repeating-Local-Global History Network for Temporal Knowledge Graph Reasoning

### v1

@@ Abstract
Temporal Knowledge Graph (TKG) reasoning is based on historical information to predict the future. Therefore, parsing and mining historical information is key to predicting the future. Most existing methods fail to concurrently address and comprehend historical information from both global and local perspectives. Neglecting the global view might result in overlooking macroscopic trends and patterns, while ignoring the local view can lead to missing critical detailed information. Additionally, some methods do not focus on learning from high-frequency repeating events, which means they may not fully grasp frequently occurring historical events. To this end, we propose the Repetitive-Local-Global History Network(RLGNet). We utilize a global history encoder to capture the overarching nature of historical information. Subsequently, the local history encoder provides information related to the query timestamp. Finally, we employ the repeating history encoder to identify and learn from frequently occurring historical events. In the evaluation on six benchmark datasets, our approach generally outperforms existing TKG reasoning models in multi-step and single-step reasoning tasks.
@@ 1 Introduction
TKG is a structured yet highly complex knowledge system. In TKG, each fact is represented by a quadruple, including subject, relation, object, and timestamp, thereby providing detailed and accurate temporal information. TKG allow us to understand the relationships between entities and capture the changes in these relationships via timestamps.
TKG extrapolation aims to predict new facts at a future timestep based on the ground truth facts within the known timeframe. Extrapolation is further classified into two types: (1) Single-step reasoning predicts future events using known ground truth facts from past timestamps within the test period. (2) Multi-step reasoning forecasts without relying on ground truth facts from the test period. By analyzing past events and identifying factors that influence event progression, we can infer the probabilities of emergent future facts. However, since future events contain unknown factors, TKG extrapolation is a challenging problem(?).
Figure 1: A figure to explain repeating, local, and global facts. ”Frequency” represents the number of occurrences of that type of event, ”Interval” represents the time difference from the last occurrence of that type of event to the time of query. The red arrows represent the facts most relevant to the query.
Predicting the future is a task that requires a deep analysis of historical data, the key to which is to discern possible patterns and trends from it. In previous work, it’s generally believed that the facts closer to the query time are more important. However, this method may cause us to overlook some facts that have the potential to impact future trends, because these facts are not necessarily always close to the query time. As shown in Figure 1, (United States, Negotiations, ?) does not appear in the local history. According to historic recurrence(?), historical facts could have repeated or even cyclical patterns. As illustrated in the same Figure 1, (United States, Negotiations, North Korea) has occurred 95 times. Therefore, we need to consider the repetitiveness, locality, and globality of historical information. Moreover, some work has shown that not every fact in the TKG is crucially important for prediction. To improve the accuracy of prediction, we need to selectively pay attention to the facts that are most relevant to the query, as the fact indicated by the red arrow in Figure 1. Therefore, finding a method that is based on the query and simultaneously considers repeating, local and global historical information is of paramount importance for extrapolation in TKG.
To address this challenge, we design and propose a model called Repeating-Local-Global History Network (RLGNet). In this model, we design repeating, local, and global history encoders respectively based on the query-oriented method. In the local history encoder, the model views the TKG as a Knowledge Graph (KG) sequence and models the facts in adjacent timestamps to obtain local historical information. In the global history encoder, the model aggregates all facts using an attention mechanism to gather global historical information. Finally, building on the foundation of local and global historical encoders, we use a repeating history encoder to further increase the scores of repetitive events. Overall, this paper makes the following contributions:
@@
    We designed three different levels of encoders to deal with the repetitive, local, and global historical information respectively.
@@
    We explored the contributions of repeating, global and local historical information to predictive performance in two different extrapolation settings.
@@
    We conducted extensive experiments on six public TKG datasets, proving the model’s effectiveness in entity prediction.
@@ 2 Related Work
Static KG Reasoning: Recent years have seen great interest and research in static Knowledge Graph (KG) reasoning models.(??) These models include distance-based models, such as TransE(?) and TransH(?), which determine the likelihood of facts by measuring the distances between entities. Another category is semantic-matching-based models, like DistMult(?) and RESCAL(?). There are also models based on Convolutional Neural Networks (CNNs), such as ConvE(?) and Conv-TransE(?), which represent entities and relations using matrices processed by convolutional kernels. Similarly, Graph Convolutional Network (GCN) models, including R-GCN(?) and VR-GCN(?), stand out for their capacity to integrate graph structures with node features. However, these models focus on static KG, and their predictive capabilities for future events are limited.
Temporal KG Reasoning: TKG reasoning has two settings: extrapolation and interpolation. The interpolation setting aims at predicting missing historical facts rather than future events. In contrast, this paper focuses on reasoning within the extrapolation setting to forecast future facts using historical data. Know-Evolve(?) and DyREP(?) model the occurrence of facts in TKG using a temporal point process. Glean(?) enriches factual features by utilizing unstructured text information. CyGNet(?) captures and understands historical trends and patterns through a replicating generative mechanism. CENET(?) distinguishes events as historical and non-historical, using this classification for contrastive learning. Some models, such as RE-GCN(?), utilize Graph Convolutional Networks (GCNs) to simulate the evolutionary process of knowledge graphs, thereby capturing and learning the dynamic properties of entities and relations. TiRGN(?) uses a local-global historical approach for reasoning. CEN(?) and CluSTeR(?) utilize reinforcement learning to capture and learn the connections between entities and relationships in the knowledge graph. TLogic(?) utilizes temporal logic rules to constrain the predicted paths for queries. TANGO(?) utilizes neural ordinary differential equations to model the structural information of each entity. However, these models often focus on partial aspects of historical information without combining query data with historical details. Consequently, they can’t accurately capture recurring, local, and global facts.
@@ 3 Preliminaries
TKG represent events temporally as snapshots. Let $\mathcal{G}_{raw}$ be the sequence of snapshot graphs $\mathcal{G}_{raw}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}\}$. Each fact in the snapshot is a quadruple $(s,r,o,t)$, where $s,o\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{T}$. It represents a relationship $r$ between subject entity $s$ and object entity $o$ at time $t$. For each quadruple $(s,r,o,t)$,an inverse relation quadruple$(o,r^{-1},s,t)$ is often added to the dataset. The entity prediction task aims to predict the missing entity in a query $q=(s_{q},r_{q},?,t_{q})$. We denote all entity embeddings in dimension $d$ with $H\in\mathbb{R}^{|\mathcal{E}|\times d}$, and relation embeddings similarly with $R\in\mathbb{R}^{|\mathcal{R}|\times d}$. In each encoder, the initial weights of H and R are randomly generated. The specific entity and relation are represented as $h\in\mathbb{R}^{d}$and $r\in\mathbb{R}^{d},$ respectively.
Given the query tuple $q=(s_{q},r_{q},?,t_{q})$, we define the candidate entity set at time $t$ as $\mathcal{C}_{q}^{t}=\{o|(s_{q},o_{q},o,t)\in\mathcal{G}_{t}\}$. The candidate entity set over a time range $t_{1}$ to $t_{2}$ is defined as $\mathcal{C}_{q}^{t_{1}:t_{2}}$, a union of all candidate sets over the range.
@@
Since not all candidate entities greatly affect the prediction, we only keep the top $k$ most frequent entities, denoted as $C_{q,top_{k}}^{1:t_{q}-1}$. Candidates in the set usually represent repeating facts, we construct $\mathcal{G}_{rep}$, containing all tuples in the period where o is a top $k$ candidate before $t-1$. $\mathcal{G}_{rep}$ can be formalized as:
$\mathcal{G}_{rep}=\{q=(s_{q},r_{q},o_{q},t_{q})\in\mathcal{G}_{raw}|o\in\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}\}$ (2)
$\mathcal{G}_{rep}$ essentially extracts prominent, repetitive events from historical data.
Figure 2: The overall architecture of RLGNet. The upper half is the local history encoder which learns local historical information through KG sequences of adjacent timestamps. The lower half comprises the repeating and global history encoders, learning repetitive and global historical information respectively by statistically querying candidate entities. The content within the dashed box represents the entity most relevant to the query. The left side sums the scores of the three parts to get the final prediction score.
@@ 4 Model Overview
The overall framework of the model is illustrated in the Figure 2. This model can be divided into three core submodules: the local history encoder, the global history encoder and the repeating history encoder. The local history encoder captures local facts by exploring the structural features and historical dependencies the sequence of KG at adjacent timestamps. The global history encoder captures global facts by examining the relevant facts across all previous timestamps and capturing entities and relationships that may not be evident at adjacent timestamps. Furthermore, these two encoders use a scoring decoder to evaluate the facts. The primary task of the repeating history encoder is to enhance its ability to predict repetitive facts by encoding frequently occurring historical events. All encoders utilize time vector to encode time and frequency. By utilizing these three encoders, the model is capable of making accurate predictions in response to queries.
4.1 Scoring Decoder
Inspired by the work of ConvE, we designed this scoring decoder and used it in the next two decoders. First, the model reshapes a one-dimensional input into a two-dimensional matrix. Then, it processes this matrix with a CNN. After that, it flattens the resulting data and outputs the embeddings for the target entity through a fully connected layer.
$\varphi(H,V_{1}\dots V_{k})=\sigma(ConvE(V_{1},\dots,V_{k}))\cdot H$ (3)
$\sigma$ represents the ReLu function. $V_{i}(1\leq i\leq k)$are input vectors, $\varphi(H,V_{1},\dots,V_{k})\in\mathbb{R}^{|\mathcal{E}|}$ represents predicted entity scores.
4.2 Time Vector
In terms of time, the module employs two types of temporal encoding vectors, periodic(?) and aperiodic, to represent time.
$\textbf{v}_{1}(t)&=\cos(W_{1}^{t}t+b_{1})\\ \textbf{v}_{2}(t)&=\tanh(W_{2}^{t}t+b_{2})\end{aligned}\right.$ (4)
Then we concatenate the two vectors to obtain the final vector representation.
$\textbf{V}(t)=[\textbf{v}_{1}(t);\textbf{v}_{2}(t)]$ (5)
$W_{1}^{t},W_{2}^{t},b_{1},b_{2}$ are learnable parameters and the vectors $\textbf{v}_{1}(t)$ and $\textbf{v}_{2}(t)$ have the same dimensionality, [;] denotes the vector concatenation operation, $\textbf{V}(t)$ is the time vector.
4.3 Local History Encoder
@@
This module captures local facts by focusing on adjacent history. For each query $q=(s_{q},r_{q},?,t_{q})$, the module pays attention to the $m$ timestamp subgraph $\{\mathcal{G}_{t_{q}-m},\dots,\mathcal{G}_{t_{q}-1}\}$ related to this query to obtain the structural features of the subgraph. For this purpose, we employ GCN to aggregate individual subgraphs, use Gated Recurrent Units (GRU)(?) to learn the features of subgraph evolution, and adopt an attention mechanism to integrate information from candidate entities. Firstly, we use GCN with $\omega$ layers to get the entity representations at the current moment.
$h_{o}^{t,l+1}=\sum\limits_{(s,r,o)\in\mathcal{G}_{t}}\frac{1}{|N_{o}^{t}|}W_{1}^{l}\phi(h_{i}^{t,l},r_{t})+W_{2}^{l}h_{o}^{l}$ (6)
$\phi(h_{i}^{t,l},r_{t})=[h_{i}^{t,l};r_{t};h_{i}^{t,l}+r_{t};h_{i}^{t,l}\cdot r_{t}]$ (7)
@@
The symbol [$\cdot$] represents the Hadamard product. $N_{o}^{t}$ represents neighbors of node $o$ at time $t$. The embedded representation of node $o$ at the $l$-th layer is denoted as $h_{o}^{t,l}\in\mathbb{R}^{d}$. The aggregation and self-loop parameters of the $l$-th layer are shown as $W_{1}^{l}$ and $W_{2}^{l}$. The edge relation type is presented as $r_{t}$. Specifically, when $\omega=0$, i.e., $h_{o}^{t,0}=\sum_{i\in N_{o}^{t}}\frac{1}{|N_{o}^{t}|}h_{i}^{t}$. Entity representations are combined with their temporal representations and then predicted for the next moment with GRU:
$H_{t+1}=GRU([H_{t};\textbf{V}_{1}(t_{q}-t)],H^{{}^{\prime}})$ (8)
@@
$\textbf{V}_{1}(t_{q}-t)\in\mathbb{R}^{d_{2}}$ is the time vector. $H^{{}^{\prime}}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the final GCN layer output. The attention mechanism aggregates candidate entities over a history of length $m$. Node aggregation for query $q$ is:
$C_{q}^{t}=\sum\limits_{i\in C_{q}^{t}}\frac{1}{|C_{q}^{t}|}h_{i}^{{}^{\prime}}$ (9)
@@
$C_{q}^{t}$ represents query $q$’s candidate entity representation at time $t$. Next, the module calculates attention weights $a_{q}^{t}$ for the query at time $t$:
$a_{q}^{t}=W_{2}\sigma(W_{1}[h_{q}^{{}^{\prime}};r_{q};\textbf{V}_{1}(t_{q}-t);C_{q}^{t}])$ (10)
@@
$W_{1}$, and $W_{2}$ are learnable parameters. After calculating attention weights, these are used to aggregate candidate entities for query q:
$C_{q}^{{}^{\prime}}=\sum_{i=1}^{m}\frac{\exp{(a_{q}^{i})}\cdot C_{q}^{i}}{\sum_{j=1}^{m}\exp{(a_{q}^{j})}}$ (11)
The candidate entity representation $C_{q}^{{}^{\prime}}$ for query $q$. The module then calculates the entity score for query $q$’s local history encoder:
$score_{loc}^{q}=\varphi_{loc}(H^{{}^{\prime}},h_{q}^{{}^{\prime}},r_{q},C_{q}^{{}^{\prime}})$ (12)
Eq (12) can be viewed as an instance of Eq (3).
4.4 Global History Encoder

### v2

@@ Abstract
Temporal Knowledge Graph (TKG) reasoning involves predicting future events based on historical information. However, due to the unpredictability of future events, this task is highly challenging. To address this issue, we propose a multi-scale hybrid architecture model based on ensemble learning, called RLGNet (Repeating-Local-Global History Network). Inspired by the application of multi-scale information in other fields, we introduce the concept of multi-scale information into TKG reasoning. Specifically, RLGNet captures and integrates different levels of historical information by combining modules that process information at various scales. The model comprises three modules: the Repeating History Module focuses on identifying repetitive patterns and trends in historical data, the Local History Module captures short-term changes and details, and the Global History Module provides a macro perspective on long-term changes. Additionally, to address the limitations of previous single-architecture models in generalizing across single-step and multi-step reasoning tasks, we adopted architectures based on Recurrent Neural Networks (RNN) and Multi-Layer Perceptrons (MLP) for the Local and Global History Modules, respectively. This hybrid architecture design enables the model to complement both multi-step and single-step reasoning capabilities. Finally, to address the issue of noise in TKGs, we adopt an ensemble learning strategy, combining the predictions of the three modules to reduce the impact of noise on the final prediction results. In the evaluation on six benchmark datasets, our approach generally outperforms existing TKG reasoning models in multi-step and single-step reasoning tasks.
Figure 1: The figure illustrates the repeating, local, and global perspectives. The red timestamp indicates that an fact identical to the queried fact occurred at that specific moment in time.
@@ 1 Introduction
A Temporal Knowledge Graph (TKG) is a structured yet highly complex knowledge system. In TKG, each fact is represented by a quadruple consisting of subject, relation, object, and timestamp. This paper primarily discusses the problem of extrapolation in TKG. TKG allows us to understand the relationships between entities and capture changes in these relationships over time. The goal of TKG extrapolation is to predict new facts at future time steps based on known facts. Extrapolation can be further divided into single-step and multi-step types. Single-step reasoning involves predicting facts that will occur at the current time step (time $t$) based on known past facts. Multi-step reasoning requires sequentially inferring facts at multiple future time steps, such as $t$, $t+1$, $t+2$, and so on. However, the inclusion of unknown factors in future events poses significant challenges for TKG extrapolation (?).
Recent studies indicate that some facts in TKG exhibit traceable patterns, hence simulating the evolution of these facts can aid in predicting them. For instance, models based on RNN architecture are widely used (?). Additionally, there are models based on MLP (?), where time is often vectorized and input as a feature to capture the temporal characteristics of facts. To achieve more accurate predictions, some models cleverly design methods to fuse local and global historical information, allowing the model to have both global and local perspectives. The global perspective helps the model better generalize the entire dataset’s characteristics, while the local perspective assists in extracting recent fact information. Furthermore, some studies identify closely related entities as candidate answers, using them to constrain the prediction range and improve accuracy.
Beyond these designs, we further observe that a portion of facts in TKG are closely related to the past, as evidenced by the significant proportion of facts that repeat multiple times in TKG. Thus, when predicting the future, it is reasonable to assume that these facts will reoccur. Conversely, some facts in TKG only occur at specific times, contributing to substantial noise in TKG datasets. This noise causes models to fit these time-specific facts during training, impairing their ability to predict future events. In TKG multi-step reasoning tasks, the absence of facts over a period can cause RNN-based models to evolve incorrect facts. Over time, the influence of past facts diminishes, and MLP-based models tend to assign similar weights to facts at different times, affecting their performance in single-step reasoning. Just as RNN and MLP models each have their strengths in single-step and multi-step reasoning tasks, they also have advantages in capturing historical information at different scales. MLP-based models are well-suited for capturing patterns and relationships across the entire dataset because they process each input independently, without temporal bias, allowing them to generalize long-term trends and patterns. On the other hand, RNN-based models are ideal for capturing local historical information, as they excel at handling short sequences of data and can effectively capture short-term changes in the most recent time steps.
Based on these observations, we designed a model called the Repeating-Local-Global History Network (RLGNet). Specifically, RLGNet draws inspiration from ensemble learning by designing three relatively independent modules: Global History Module, Local History Module, and Repeating History Module. As shown in Figure 1, these three modules capture information at different scales. During training, RLGNet first learns the Local and Global History Modules in parallel, while the Repeating History Module learns serially on the results of the Global and Local History Modules to further reinforce the scores of repeating facts. This ensemble learning strategy can reduce the sensitivity of individual models to noisy data in TKG, better capture the true patterns in the data, and reduce the risk of overfitting. Due to the advantages of RNN-based and MLP-based models in single-step and multi-step reasoning tasks, as well as their effectiveness in handling historical information at different scales, RLGNet adopts an RNN-based structure for the Local History Module and an MLP-based structure for the Global History Module. This hybrid design fully leverages the complementary strengths of both approaches. Finally, to further constrain the answer range, we integrate candidate entities’ information into the three learning processes. Our multi-scale hybrid architecture design enables RLGNet to achieve state-of-the-art performance in most cases for both multi-step and single-step reasoning tasks. Our contributions are summarized as follows:
@@
    Unlike previous work, we designed an ensemble learning-based model that processes different historical information from a multi-scale perspective. Specifically, we use Repeating, Local, and Global History Modules to capture their corresponding historical information, and these modules collaborate to enhance the overall performance of the model.
@@
    We use RNN and MLP architectures in the Local History Module and Global History Module, respectively, and integrate their scores to achieve complementary capabilities between the two modules. This hybrid architecture model addresses the limitation of previous works, where a single architecture could only achieve good results in either multi-step or single-step tasks, improving the model’s performance in multi-task scenarios.
@@
    We proposed a multi-scale hybrid architecture model based on an ensemble learning strategy, which demonstrated impressive performance across six public TKG datasets.
@@ 2 Related Work
Static KG Reasoning. Recent years have seen great interest and research in static Knowledge Graph (KG) reasoning models.(??) These models include distance-based models, such as TransE(?) and TransH(?), which determine the likelihood of facts by measuring the distances between entities. Another category is semantic-matching-based models, like DistMult(?) and RESCAL(?). There are also models based on Convolutional Neural Networks (CNNs), such as ConvE(?) and Conv-TransE(?), which represent entities and relations using matrices processed by convolutional kernels. Similarly, Graph Convolutional Network (GCN) models, including R-GCN(?) and VR-GCN(?), stand out for their capacity to integrate graph structures with node features. However, these models focus on static KG, and their predictive capabilities for future events are limited.
Temporal KG Reasoning. TKG reasoning has two settings: extrapolation and interpolation. The interpolation setting aims at predicting missing historical facts rather than future events. In contrast, this paper focuses on reasoning within the extrapolation setting to forecast future facts using historical data. Know-Evolve(?) and DyREP(?) model the occurrence of facts in TKG using a temporal point process. Glean(?) enriches factual features by utilizing unstructured text information. CyGNet(?) captures and understands historical trends and patterns through a replicating generative mechanism. CENET(?) distinguishes events as historical and non-historical, using this classification for contrastive learning. Some models, such as RE-GCN(?), utilize Graph Convolutional Networks (GCNs) to simulate the evolutionary process of knowledge graphs, thereby capturing and learning the dynamic properties of entities and relations. TiRGN(?) uses a local-global historical approach for reasoning. CEN(?) and CluSTeR(?) utilize reinforcement learning to capture and learn the connections between entities and relationships in the knowledge graph. TLogic(?) utilizes temporal logic rules to constrain the predicted paths for queries. TANGO(?) utilizes neural ordinary differential equations to model the structural information of each entity. However, these models often focus on partial aspects of historical information without combining query data with historical details. Consequently, they can’t accurately capture recurring, local, and global facts.
@@ 3 Preliminaries
TKGs represent events temporally as snapshots. Let $\mathcal{G}_{raw}$ be the sequence of snapshot graphs $\mathcal{G}_{raw}=\{\mathcal{G}_{raw}^{1},\dots,\mathcal{G}_{raw}^{|\mathcal{T}|}\}$. Each fact in the snapshot is a quadruple $(s,r,o,t)$, where $s,o\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{T}$. It represents a relationship $r$ between subject entity $s$ and object entity $o$ at time $t$. $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{T}$ represent the sets of entities, relations, and timestamps, respectively. In this paper, our reasoning task is to predict missing entities, specifically the subject entity $s$ and the object entity $o$. To simplify the problem, we convert the task of predicting the object entity $o$ into predicting the subject entity $s$, represented as $(o,r^{-1},?,t)$. Therefore, for each quadruple $(s,r,o,t)$, we add a reverse relation quadruple $(o,r^{-1},s,t)$ to the dataset.
Given the query tuple $q=(s_{q},r_{q},?,t_{q})$, we define the candidate entity set at time $t$ as $\mathcal{C}_{q}^{t}=\{o|(s_{q},r_{q},o,t)\in\mathcal{G}^{t}_{raw}\}$. The candidate entity set over a time range $t_{1}$ to $t_{2}$ is defined as $\mathcal{C}_{q}^{t_{1}:t_{2}}$, a union of all candidate sets over the range:
@@
Since not all candidate entities greatly affect the prediction, we only keep the top $top_{k}$ most frequent entities, denoted as $\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$. To represent repeating facts, we need to select those facts that have occurred more than once in history and define repeating facts that occurred at time $t$ as $\mathcal{G}_{rep}^{t}$. $\mathcal{G}_{rep}^{t}$ can be formalized as:
$\mathcal{G}_{rep}^{t}=\{q^{\prime}|q^{\prime}=(s,r,o,t)\in\mathcal{G}^{t}_{raw},o\in\mathcal{C}_{q^{\prime},top_{k}}^{1:t-1}\}$ (2)
Similar to $\mathcal{G}_{raw}$, we obtain $\mathcal{G}_{rep}=\{\mathcal{G}_{rep}^{1},\dots,\mathcal{G}_{rep}^{|\mathcal{T}|}\}$.
Figure 2: The upper half represents the Local History Module, which learns local historical information through KG sequences of adjacent timestamps. The lower half comprises the Repeating and Global History Modules, learning repetitive and global historical information respectively by statistically querying candidate entities. The entities within the dashed box are candidate entities. The left side weights and sums the scores of the three modules to obtain the final prediction score.
@@ 4 Model Overview
The overall framework of the model is illustrated in Figure 2. This model can be divided into three core submodules: the Local History Module, the Global History Module, and the Repeating History Module. The Local and Global History Modules are used to capture historical information at local and global scales, respectively, with the Local History Module based on an RNN architecture and the Global History Module based on an MLP architecture. The Repeating History Module is used to further enhance the scores of repeating facts. To handle the frequency of time and entity occurrences, we use a Numerical Embedding function to convert them into embeddings. To further constrain the answer range, we apply the candidate entity design to all three modules. The model is designed based on the concept of ensemble learning, so the repeating, global, and local modules are trained independently, and their prediction scores are combined through weighted summation to produce the final result.
4.1 Numerical Embedding
The Numerical Embedding module is used to convert time and entity occurrence frequency into embeddings. The cosine function can capture periodic patterns in time and frequency, while the tanh function can provide smoother transitions for phenomena with rapidly changing frequencies or sudden changes in time.
$\textbf{v}_{1}(x)&=\cos(W_{1}^{N}x+b_{1}^{N})\\ \textbf{v}_{2}(x)&=\tanh(W_{2}^{N}x+b_{2}^{N})\end{aligned}\right.$ (3)
Then we concatenate the two embeddings to obtain the final embedding.
$\textbf{V}(t)=[\textbf{v}_{1}(x);\textbf{v}_{2}(x)]$ (4)
Here, $W_{1}^{N},W_{2}^{N},b_{1}^{N},b_{2}^{N}$ are learnable parameters and the embeddings $\textbf{v}_{1}(x)$ and $\textbf{v}_{2}(x)$ have the same dimensionality. The notation [;] denotes the embedding concatenation operation, and $\textbf{V}(x)$ represents the Numerical Embedding. Combining the embeddings generated by these two modules with different characteristics allows for the integration of both periodic and nonlinear information, providing a more comprehensive representation.
4.2 Local History Module
@@
This module captures local facts by focusing on adjacent history. For each query $q=(s_{q},r_{q},?,t_{q})$, the module pays attention to the $m$ timestamp subgraph $\{\mathcal{G}^{t_{q}-m}_{raw},\dots,\mathcal{G}^{t_{q}-1}_{raw}\}$ related to this query to obtain the structural features of the subgraph. Here, $m$ is a hyperparameter. To achieve this, we employ a GCN to aggregate individual subgraphs, use Gated Recurrent Units (GRU) (?) to learn the features of subgraph evolution, and adopt an attention mechanism to integrate information from candidate entities. Firstly, we use a GCN with $\omega$ layers to get the entity embeddings at the current moment:
$h_{o}^{t,l+1}=\sum\limits_{(s,r,o,t)\in\mathcal{G}^{t}_{raw}}\frac{1}{|N_{o}^{t}|}W_{1}^{l}\phi(h_{s}^{t,l},r^{L})+W_{2}^{l}h_{o}^{l}$ (5)
$\phi(h_{s}^{t,l},r^{L})=[h_{s}^{t,l};r^{L};h_{s}^{t,l}+r^{L};h_{s}^{t,l}\cdot r^{L}]$ (6)
@@
Here, the symbol [$\cdot$] represents the Hadamard product. $N_{o}^{t}$ represents the neighbors of node $o$ at time $t$. The embedded of node $o$ and $s$ at the $l$-th layer is denoted as $h_{o}^{t,l}$ and $h_{s}^{t,l}$. The aggregation and self-loop parameters of the $l$-th layer are denoted as $W_{1}^{l}$ and $W_{2}^{l}$. $r^{L}$ is the initial embedding of the relation in the Local History Module. When $\omega=0$,$h_{o}^{t,0}=\sum_{i\in N_{o}^{t}}\frac{1}{|N_{o}^{t}|}h_{i}^{t}$. If $t=t_{q}-m$ then $h^{t_{q}-m}_{o}$ represents the initial embedding of the entity $o$ in the Local History Module. Entity embeddings are combined with their temporal embeddings and then predicted for the next moment with a GRU:
$H_{t+1}=GRU([H_{t};\textbf{V}^{L}(t_{q}-t)],H^{{}^{\prime}}_{t})$ (7)
@@
Here, $\textbf{V}^{L}$ is Numerical Embeddings. $H^{{}^{\prime}}_{t}$ is the output of the last layer of the GCN for all entities at time $t$, and $H_{t}$ is the embedding representation of all entities at time $t$. The attention mechanism aggregates candidate entities over a history of length $m$. Node aggregation for query $q$ is:
$C_{q}^{t}=\sum\limits_{i\in\mathcal{C}_{q}^{t}}\frac{1}{|\mathcal{C}_{q}^{t}|}h_{i}^{t}$ (8)
@@
Here, $C_{q}^{t}$ represents query $q$’s candidate entity embedding at time $t$. $h_{i}^{t}$ represents the embedding of candidate entity $i$ in $H_{t}$. Next, the module uses a 2-layer MLP to calculate the attention weights $a_{q}^{t}$ for the query at time $t$:
$a_{q}^{t}=MLP([h_{q}^{t};r_{q}^{L};\textbf{V}^{L}(t_{q}-t);C_{q}^{t}])$ (9)
@@
Here, $r_{q}^{L}$ is the initial embedding of the query relation in the Local History Module. $h_{q}^{t}$ represents the embedding of the query entity in $H_{t}$. After calculating attention weights, these are used to aggregate candidate entities for query q:
$C_{q}^{{}^{\prime}}=\sum_{i=t_{q}-m}^{t_{q}-1}\frac{\exp{(a_{q}^{i})}\cdot C_{q}^{i}}{\sum_{j=t_{q}-m}^{t_{q}}\exp{(a_{q}^{j})}}$ (10)
This results in the candidate entity embedding $C_{q}^{{}^{\prime}}$ for query $q$. The Local History Module uses ConvTransE as the scoring function. To better adapt to the task, we additionally concatenate query-related information embeddings to the original entity and relation embeddings in ConvTransE. These concatenated embeddings are then processed through a one-dimensional convolution layer followed by a fully connected layer to obtain the scores:
$score_{loc}^{q}=H_{t_{q}}\cdot\phi_{L}(h_{q}^{t_{q}},r_{q}^{L},C_{q}^{{}^{\prime}})$ (11)
Here, $\phi_{L}$ represents the score decoder based on ConvTransE in the Local History Module.
4.3 Global History Module